# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v5

@@ 1 Introduction
Table 1: Comparison of ITE Estimation Approaches (Hypertension Medication Example). $ITE=Y(1)-Y(0)$ represents individual treatment effect, $\tau$ is a clinical threshold (e.g., 15 mmHg reduction).
@@
[…]
Possible Theoretical Guarantee Consistent for CATE Marginal coverage Probabilistic and quantile calibration
Reliable distributional estimation for ITEs must address three key challenges: (1e) Confounding robustness to covariate shift via propensity score adjustment; (2) Universal asymptotic convergence to the data-generating ITE distribution; and (3) Finite sample calibration even for small $n$. As can be seen in Table 2, existing distributional approaches struggle mostly with the third challenge to provide calibrated predictive distributions for counterfactuals and ITEs. Additionally, none of the existing solutions is model agnostic. Bayesian methods (BART (Chipman et al., 2010; Hill, 2011), CMGP (Alaa and van der Schaar, 2017), DKLITE (Zhang et al., 2020)) and some deep learning approaches (CEVAE (Louizos et al., 2017)) impose restrictive parametric assumptions limiting universality. While other neural network approaches (GANITE (Yoon et al., 2018), NOFLITE (Vanderschueren et al., 2023), DiffPO (Ma et al., 2024), FCCN (Zhou et al., 2022)) allow for the modeling of more flexible predictive distributions, they often produce poorly calibrated distributions and lack finite-sample guarantees of probabilistic calibration. Conformal prediction (CP) approaches (WCP (Lei and Candès, 2021), CM-learners (Alaa et al., 2023)) advance interval estimation but cannot quantify threshold probabilities or integrate utility functions. These limitations persist because counterfactual inference fundamentally involves unobserved outcomes, making it a challenge to estimate predictive distributions while balancing robustness, calibration, and computational tractability.
@@
Approach Methods S-, T-, or Direct- learner Confounding Universal Finite sample Model agnostic
@@ 3.  Empirical robustness
Code and datasets are available at https://github.com/predict-idlab/cct-cmc.
@@ 2 Problem definition
$\pi(x)=\mathbb{P}(W=1|X=x),\quad W\sim\text{Bernoulli}(\pi(x)),\quad X\sim\Lambda$
$Y(w)=\mu_{w}(X)+\varepsilon(w),\quad\mu_{w}(x)=\mathbb{E}(Y|X=x,W=w),\quad Y=Y(0)(1-W)+Y(1)W$
$\tau(x)=\mathbb{E}(Y(1)-Y(0)|X=x),\quad ITE=Y(1)-Y(0)=\tau(X)+\varepsilon_{a,ITE}$
@@ 3 Conformal predictive systems
Vovk et al. (2019, 2020) propose conformal predictive systems (CPS), an extension of CP (Appendix A.1) that allows deriving predictive distributions and enjoys provable properties of validity, in the sense of probabilistic calibration, under exchangeability (Vovk et al., 2019). CPS produces conformal predictive distributions as p-values arranged into a probability distribution function Vovk et al. (2019), created with the help of conformity scores. Vovk et al. (2019) define a CPS as a function that is both a smooth conformal transducer and a randomized predictive system (RPS). The conformal transducer is another way to represent a conformal predictor, i.e., a confidence transducer that can be seen as a function that maps each input sequence into a sequence of p-values (Vovk et al., 2022). Formally, we define the full smoothed conformal transducer, defined by conformity score $s$, as
$Q(X_{n+1},y,\phi;Z_{1:n}):=\sum_{i=1}^{n+1}[S_{i}^{y}<S_{n+1}^{y}]\frac{1}{n+1}+\sum_{i=1}^{n+1}[S_{i}^{y}=S_{n+1}^{y}]\frac{\phi}{n+1}$
@@
A smoothed conformal transducer is an RPS if and only if the exchangeability assumption is satisfied and if it is defined by a monotonic conformity score (Vovk et al., 2019). CPS has been adapted to become more computationally efficient, i.e., split conformal predictive systems (SCPS) (Vovk et al., 2020), by building upon inductive CP (ICP). Here, the p-values are created by a split conformity score function that is an element of a family of measurable functions $s:\mathcal{Z}\times\mathcal{Z}^{m}\rightarrow\mathbb{R}\cup\{-\infty,\infty\}$, where $m$ is the length of the training sequence. This score needs to be isotonic and balanced. Examples of such a measure are the residual error $y-\hat{f}(x)$ and normalized residual error $\frac{y-\hat{f}(x)}{\hat{u}(x)}$, where $\hat{f}(\cdot)$ represent the fitted regression model and $\hat{u}(\cdot)$ a 1D-uncertainty estimate. The SCPS is a function that is both a split-smoothed conformal transducer and an RPS and follows a similar form as in Eq. 3, however, only leveraging conformity scores of the calibration set. Vovk et al. (2020) proved that the split smoothed conformal transducer is an RPS if the exchangeability assumption is satisfied and the conformity measure is isotonic and balanced. In this work, we also introduce a new conformity score for CPS, namely, the probability integral transform (PIT)-values $s=\hat{F}_{X_{i}}(Y_{i})$ of the distribution prediction $\hat{F}_{X_{i}}$. For this score, you need access to the CDF estimate and thus require a distributional regression approach as a base-learner. Note that Chernozhukov et al. (2021) introduced a similar score in their distributional CP approach.
@@ 3.1 Conformal predictive systems under covariate shift
Assume that $Z_{i}=(X_{i},Y_{i})\in\mathcal{X}\times\mathbb{R}$, $i=1,...,n$ are produced independently from $\mathcal{P}=\mathcal{P}_{X}\times\mathcal{P}_{Y|X}$; $Z_{n+1}=(X_{n+1},Y_{n+1})\in\mathcal{X}\times\mathbb{R}$, is independently drawn from $\tilde{\mathcal{P}}=\tilde{\mathcal{P}}_{X}\times\mathcal{P}_{Y|X}$; $\tilde{\mathcal{P}}_{X}$ is absolutely continuous with respect to $\mathcal{P}_{X}$; likelihood ratio $w(x)=\frac{d\tilde{\mathcal{P}}_{X}}{d\mathcal{P}_{X}}$ representing the covariate shift, random number $\phi\sim Uniform(0,1)$; $Z_{1:n}$, $Z_{n+1}$, and $\phi$ to be independent. Then, the distribution of the weighted conformal transducer, defined by
$Q(X_{n+1},y,\phi,w(x);Z_{1:n}):=\sum_{i=1}^{n+1}[S_{i}^{y}<S_{n+1}^{y}]p_{i}^{w}+\sum_{i=1}^{n+1}[S_{i}^{y}=S_{n+1}^{y}]p_{i}^{w}\phi$ (1)
@@
$\forall\alpha\in[0,1]:\mathbb{P}_{Z_{1:n}\sim\mathcal{P},(X_{n+1},Y_{n+1})\sim\tilde{\mathcal{P}}}\{Q(X_{n+1},Y_{n+1},\phi,w(x);Z_{1:n})\leq\alpha\}=\alpha$ (2)
@@ 4 Conformal convolution and conformal Monte Carlo meta-learner
3:  $Z^{W=0}_{1:n_{0}}:=\{(Y_{i},X_{i})|W_{i}=0\}_{i=1}^{n}$ where $n_{0}$ the number observation under control.
4:  $Z^{W=1}_{1:n_{1}}:=\{(Y_{i},X_{i})|W_{i}=1\}_{i=1}^{n}$ where $n_{1}$ the number observation under treatment.
@@
3:  $Z_{train,1:n_{train}},Z_{cal,1:n_{cal}}=\textbf{RandomSplit}(Z_{1:n})$
4:  $Z^{W=0}_{train,1:n_{0}}:=\{(Y_{train,i},X_{train,i})|W_{train,i}=0\}_{i=1}^{n_{train}}$ where $n_{0}$ the number observation in the training split under control.
5:  $Z^{W=1}_{train,1:n_{1}}:=\{(Y_{train,i},X_{train,i})|W_{train,i}=1\}_{i=1}^{n_{train}}$ where $n_{1}$ the number observation in the training split under treatment.
@@
7:  $ITE_{MC,1:n_{ITE}},X_{MC,1:n_{ITE}}=\textbf{MonteCarloITE}(Z_{cal,1:n_{cal}},\textbf{Q}_{Y(0)},\textbf{Q}_{Y(1)},n_{MC})$ where $n_{ITE}=n_{cal}n_{MC}.$Algorithm 3}
@@
From WCPS, we infer the potential outcomes predictive distributions, $Q_{-Y(0)}$ and $Q_{Y(1)}$, which are probabilistically calibrated. Since our primary interest often lies in ITE, defined as $ITE=Y(1)-Y(0)$, we introduce a convolution-like operation on the treatment and the negated control predictive distributions, as defined in Eq. 3. Theorem 2 states the finite-sample calibration guarantee of the CCT-learner, and the proof is given in Appendix E.2.
@@
Assume that $Z_{i}=(X_{i},W_{i},Y_{i})\in\mathcal{X}\times\{0,1\}\times\mathbb{R}$, $i=1,...,n$ and $X_{n+1}\in\mathcal{X}$, $Y_{n+1}(0),Y_{n+1}(1)\in\mathbb{R}$ are produced independently from $\mathcal{P}$, and $\phi\sim Uniform(0,1)$.
@@
The "Epsilon Problem" highlights a fundamental challenge in the predictive inference of ITE: the joint distribution of counterfactual noise terms, $\epsilon(0)$ and $\epsilon(1)$, is unidentifiable as they are never observed simultaneously. Existing approaches implicitly assume independent noise across treatments, or an identical relation, assumptions that are untestable yet critical for valid predictive distributions. We formalize that the ITE distribution is only identifiable if potential outcome marginals are independent or linked by a specified copula (Proposition 1). Violations of independence—such as correlated noise due to shared confounders—compromise distributional validity, though conservatively valid prediction intervals remain achievable. Simulations (Figure 11) empirically demonstrate coverage degradation under noise dependency. While partial identification via bounds offers an alternative, we pragmatically adopt the independent copula assumption (making a fourth assumption), aligning with formal causal models, to enable actionable uncertainty quantification despite its philosophical limitations. In Appendix D, we further elaborate on this problem.
@@
The CCT-learner’s finite-sample calibration guarantee hinges on untestable assumptions: (1) known propensity scores (typically estimated in practice); and (2) weighted exchangeability between pseudo ITE conformity scores and the true ITE $ITE_{n+1}$ (unverifiable due to unobserved counterfactuals). While these assumptions are theoretically critical, empirical results across synthetic and semi-synthetic experiments demonstrate near-perfect probabilistic calibration (e.g., dispersion of PIT values in Section 5.2 and PIT plots in Appendix F), suggesting robustness in practice. Further limitations arise from data efficiency trade-offs inherent in split CPS, as well as the sensitivity to base model’s accuracy.
@@ 5.2 Results
In our semi-synthetic experiments, we always estimate our propensity scores using logistic regression to represent a more realistic setting. Results show that we can generate probabilistically calibrated predictive distributions even when the propensity score is estimated by checking the dispersion and distribution of PIT-values (Figure 2, 8, 9).
In general, we see that most related work approaches have difficulty generating probabilistically calibrated distributions (Figure 9). We see that BART and CMGP are reasonably well calibrated when the noise distribution of the experiment is homoscedastic and normally distributed (Figure 9 for Alaa et al. (2023)). However, once the noise becomes heteroscedastic and certainly nonnormal, their calibration deteriorates (Figure 2, 9). We also observe that the CCT-learner can probabilistically calibrate the FCCN and NOFLITE approaches, often generating slightly overconfident predictive distributions.
@@
[…]
@@ A.1 Conformal prediction
$\begin{split}\hat{C}_{\alpha}&(X_{n+1})=\Bigl{\{}\\ &y\in\mathbb{R}:\frac{|i=1,...,n+1:s((X_{i},Y_{i});Z_{1:n\setminus i}\cup(X_{n+1},y))\geq s((X_{n+1},y);Z_{1:n})|}{n+1}>\alpha\\ \Bigl{\}}&\end{split}$ (6)
This procedure, termed full or transductive CP, is computationally intensive, as it requires refitting of the underlying model for each $y$ in the target domain. Therefore, a more applicable variant of full CP, called inductive or split CP (ICP) [Papadopoulos et al., 2002], offers computational efficiency and enables the integration of CP with ML algorithms, such as neural networks and tree-based models. In ICP, the training sequence $Z_{1:n}=\{(X_{1},Y_{1}),...,(X_{n},Y_{n})\}$ is portioned into a proper training sequence $\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ and a calibration sequence $\{(X_{m+1},Y_{m+1}),...,(X_{n},Y_{n})\}$. We first train a regression model $\hat{y}(x)$ using the proper training sequence. Then, based on nonconformity score $s$, we generate nonconformity scores $S_{i}$ for data points $(X_{i},Y_{i})$ where $m<i\leq n$. The nonconformity score function $s$ leverages the pre-trained regression model, e.g., the absolute error $s=|y-\hat{y}|$. These nonconformity scores together with test object $X_{n+1}$ and significance level $\alpha$ defines:

### v6

@@ 1 Introduction
Table 1: Comparison of ITE Estimation Approaches (Hypertension Medication Example). ITE=Y(1)Y(0)ITE=Y(1)-Y(0) represents individual treatment effect, $\tau$ is a clinical threshold (e.g., 15 mmHg reduction).
@@
Utility
Integration
Only computes $U(12)$ Requires worst-case assumptions Enables $\mathbb{E}[U(ITE)]=\int U(y)dF_{ITE}(y)$
Possible Theoretical
Guarantee
Consistent for CATE Marginal coverage Probabilistic and quantile calibration
Reliable distributional estimation for ITEs must address three key challenges: (1) Confounding robustness to covariate shift via propensity score adjustment; (2) Universal asymptotic convergence to the data-generating ITE distribution; and (3) Finite sample calibration even for small $n$. As can be seen in Table 2, existing distributional approaches struggle mostly with the third challenge to provide calibrated predictive distributions for counterfactuals and ITEs. Additionally, none of the existing solutions is model agnostic. Bayesian methods (BART (Chipman et al., 2010; Hill, 2011), CMGP (Alaa and van der Schaar, 2017), DKLITE (Zhang et al., 2020)) and some deep learning approaches (CEVAE (Louizos et al., 2017)) impose restrictive parametric assumptions limiting universality. While other neural network approaches (GANITE (Yoon et al., 2018), NOFLITE (Vanderschueren et al., 2023), DiffPO (Ma et al., 2024), FCCN (Zhou et al., 2022)) allow for the modeling of more flexible predictive distributions, they often produce poorly calibrated distributions and lack finite-sample guarantees of probabilistic calibration. Conformal prediction (CP) approaches (WCP (Lei and Candès, 2021), CM-learners (Alaa et al., 2023)) advance interval estimation but cannot quantify threshold probabilities or integrate utility functions. These limitations persist because counterfactual inference fundamentally involves unobserved outcomes, making it a challenge to estimate predictive distributions while balancing robustness, calibration, and computational tractability.
@@
Approach Methods
S-, T-, or Direct-
learner
Confounding
Universal
Finite
sample
Model
agnostic
@@ 3.  Empirical robustness
Code and datasets are available at https://github.com/predict-idlab/cct-cmc. Code and datasets are available in the supplementary material.
@@ 2 Problem definition
$\pi(x)=\mathbb{P}(W=1|X=x),\quad W\sim\text{Bernoulli}(\pi(x)),$
$X\sim\Lambda,\quad Y(w)=\mu_{w}(X)+\varepsilon(w),$
$\mu_{w}(x)=\mathbb{E}(Y|X=x,W=w),$
$Y=Y(0)(1-W)+Y(1)W,$
$\tau(x)=\mathbb{E}(Y(1)-Y(0)|X=x),$
$ITE=Y(1)-Y(0)=\tau(X)+\varepsilon_{a,ITE}$
@@ 3 Conformal predictive systems
Vovk et al. (2019, 2020) propose conformal predictive systems (CPS), an extension of CP (Appendix A.1) that allows deriving predictive distributions and enjoys provable properties of validity, in the sense of probabilistic calibration, under exchangeability (Vovk et al., 2019). CPS produces conformal predictive distributions as p-values arranged into a probability distribution function Vovk et al. (2019), created with the help of conformity scores. Vovk et al. (2019) define a CPS as a function that is both a smooth conformal transducer and a randomized predictive system (RPS). An RPS is a function that, given a training dataset, a test input, a candidate label $y$, and a random variable $\phi\sim Uniform(0,1)$, outputs a value in $[0,1]$ representing the cumulative probability of observing a label less than or equal to $y$. It is required to be monotonic in both $y$ and $\phi$ converges to 0 and 1 as $y\rightarrow$-$\infty$ and $y\rightarrow\infty$, and produces uniformly calibrated predictive distributions, meaning it behaves like a valid CDF and is calibrated under exchangeability. This RPS is used to formalize the theoretical framework of CPS, enabling us to prove that CPS distributional predictions have finite-sample validity. The conformal transducer is another way to represent a conformal predictor, i.e., a confidence transducer that can be seen as a function that maps each input sequence into a sequence of p-values (Vovk et al., 2022). Formally, we define the full smoothed conformal transducer, defined by conformity score $s$, as
$Q(X_{n+1},y,\phi;Z_{1:n})$ $:=\sum_{i=1}^{n+1}[S_{i}^{y}<S_{n+1}^{y}]\frac{1}{n+1}+$
$\sum_{i=1}^{n+1}[S_{i}^{y}=S_{n+1}^{y}]\frac{\phi}{n+1}$
@@
[…]
@@ 3.1 Conformal predictive systems under covariate shift
Assume that $Z_{i}=(X_{i},Y_{i})\in\mathcal{X}\times\mathbb{R}$, i=1,,ni=1,...,n are produced independently from $\mathcal{P}=\mathcal{P}_{X}\times\mathcal{P}_{Y|X}$; $Z_{n+1}=(X_{n+1},Y_{n+1})\in\mathcal{X}\times\mathbb{R}$, is independently drawn from $\tilde{\mathcal{P}}=\tilde{\mathcal{P}}_{X}\times\mathcal{P}_{Y|X}$; $\tilde{\mathcal{P}}_{X}$ is absolutely continuous with respect to $\mathcal{P}_{X}$; likelihood ratio $w(x)=\frac{d\tilde{\mathcal{P}}_{X}}{d\mathcal{P}_{X}}$ representing the covariate shift, random number $\phi\sim Uniform(0,1)$; $Z_{1:n}$, $Z_{n+1}$, and $\phi$ to be independent. Then, the distribution of the weighted conformal transducer, defined by
$\begin{split}Q(X_{n+1},y,\phi,w(x);Z_{1:n}):=\\ \sum_{i=1}^{n+1}[S_{i}^{y}<S_{n+1}^{y}]p_{i}^{w}+\sum_{i=1}^{n+1}[S_{i}^{y}=S_{n+1}^{y}]p_{i}^{w}\phi\end{split}$ (1)
@@
$\begin{split}\forall\alpha\in[0,1]:\mathbb{P}_{Z_{1:n}\sim\mathcal{P},(X_{n+1},Y_{n+1})\sim\tilde{\mathcal{P}}}\{\\ Q(X_{n+1},Y_{n+1},\phi,w(x);Z_{1:n})\leq\alpha\}=\alpha\end{split}$ (2)
@@ 4 Conformal convolution and conformal Monte Carlo meta-learner
[…]
@@
[…]
@@
[…]
@@
From WCPS, we infer the potential outcomes predictive distributions, $Q_{-Y(0)}$ and $Q_{Y(1)}$, which are probabilistically calibrated. Since our primary interest often lies in ITE, defined as ITE=Y(1)Y(0)ITE=Y(1)-Y(0), we introduce a convolution-like operation on the treatment and the negated control predictive distributions, as defined in Eq. 3. Theorem 2 states the finite-sample calibration guarantee of the CCT-learner, and the proof is given in Appendix E.2.
{theobox}
@@
Assume that $Z_{i}=(X_{i},W_{i},Y_{i})\in\mathcal{X}\times\{0,1\}\times\mathbb{R}$, i=1,,ni=1,...,n and $X_{n+1}\in\mathcal{X}$, $Y_{n+1}(0),Y_{n+1}(1)\in\mathbb{R}$ are produced independently from $\mathcal{P}$, and $\phi\sim Uniform(0,1)$.
@@
The "Epsilon Problem" highlights a fundamental challenge in the predictive inference of ITE: the joint distribution of counterfactual noise terms, $\varepsilon(0)$ and $\varepsilon(1)$, is unidentifiable as they are never observed simultaneously. Existing approaches implicitly assume independent noise across treatments, or an identical relation, assumptions that are untestable yet critical for valid predictive distributions. We formalize that the ITE distribution is only identifiable if potential outcome marginals are independent or linked by a specified copula (Proposition 1). Violations of independence—such as correlated noise due to shared confounders—compromise distributional validity, though conservatively valid prediction intervals remain achievable. Simulations (Figure 11) empirically demonstrate coverage degradation under noise dependency. While partial identification via bounds offers an alternative, we pragmatically adopt the independent copula assumption (making a fourth assumption), aligning with formal causal models, to enable actionable uncertainty quantification despite its philosophical limitations. In Appendix D, we further elaborate on this problem.
@@
The CCT-learner’s finite-sample calibration guarantee hinges on untestable assumptions: (1) known propensity scores (typically estimated in practice); and (2) weighted exchangeability between pseudo ITE conformity scores and the true ITE $ITE_{n+1}$ (unverifiable due to unobserved counterfactuals). In randomized controlled trials the propensity is known, whereas in observational studies it must be estimated, in which case the theoretical guarantee no longer strictly holds. Nonetheless, empirical results (Figures 8 and 2) suggest that the methods remain robust, and we note that other works providing coverage guarantees for counterfactual prediction intervals (Alaa et al., 2023) rely on the same assumption, and approaches (Lei and Candès, 2021) that allow an amount of mispecification in the propensity modelling guarantee conservative prediction intervals; however, a translation to predictive distributions would mean partialy unidentified distributions.
@@ 5.2 Results
In our semi-synthetic experiments, we always estimate our propensity scores using logistic regression to represent a more realistic setting. Results show that we can generate probabilistically calibrated predictive distributions even when the propensity score is estimated by checking the dispersion and distribution of PIT-values (Figure 2, 8, 8, 9).
In general, we see that most related work approaches have difficulty generating probabilistically calibrated distributions (Figure 9). We see that BART and CMGP are reasonably well calibrated when the noise distribution of the experiment is homoscedastic and normally distributed (Figure 9, Alaa experiments). However, once the noise becomes heteroscedastic and certainly nonnormal, their calibration deteriorates (Figure 8, 9). We also observe that the CCT-learner can probabilistically calibrate the FCCN and NOFLITE approaches, often generating slightly overconfident predictive distributions.
@@
The continuous ranked probability score (CRPS) measures the discrepancy between the predicted CDF and the true outcome, rewarding both calibration (statistical compatibility of probabilistic forecasts and observations) and sharpness (concentration of the distribution around the true value) (Matheson and Winkler, 1976). Mathematically defined as $\text{CRPS}(F,y)=\int(F(x)-\mathbbm{1}[x\geq y])^{2}dx$, it is a proper scoring rule as formalized by Gneiting and Raftery (2007). Lower CRPS values indicate better performance.
Table 3: Mean CRPS ranks across potential outcomes and ITE (lower is better)
Method ITEITE $Y(0)$ $Y(1)$ Avg. Rank
Alaa EDU ACIC Alaa EDU ACIC Alaa EDU ACIC
CCT-learner 1 5 1 2 6 1 1 5 1 3.07
CMC-T-learner 2 6 2 3 7 2 2 7 2 3.93
BART 5 7 6 1 8 5 3 8 5 5.00
CEVAE 8 10 7 8 9 7 9 9 7 6.93
CMGP 9 9 8 10 11 9 10 11 9 9.27
DKLITE 11 11 9 11 10 8 11 10 8 9.87
GANITE 10 8 10 9 5 10 6 6 10 8.07
NOFLITE 6 1 4 6 1 4 7 1 4 4.13
CCT-learner-NOFLITE 7 2 3 7 2 3 8 2 3 5.47
DiffPO 12 12 12 12 12 12 12 12 12 12.00
FCCN 3 4 11 5 4 11 5 4 11 6.53
CCT-learner-FCCN 4 3 5 4 3 6 4 3 6 3.73
As summarized in Table 3, the CCT- and CMC-based learners consistently rank among the top methods across ITEs and counterfactual outcomes. For the ACIC 2016 experiments, the CCT- and CMC T-learners achieve significantly lower CRPS than benchmark methods for potential outcomes and ITE distributions, demonstrating superior accuracy in capturing complex data-generating processes (Figure 2). In the synthetic experiments from Alaa et al. (2023), setup A (where treatment effects are absent), BART and CMGP perform competitively due to their S-learner structure (modeling treatment as a covariate) and alignment with homoscedastic Gaussian noise assumptions (Figure 3). However, their performance deteriorates in the semi-synthetic EDU experiment where outcomes follow an exponential noise distribution, violating their parametric assumptions (Figure 8). Here, FCCN and NOFLITE excel by flexibly modeling non-normal noise, and their integration with the CCT-learner further enhances performance by enforcing probabilistic calibration, resulting in even lower CRPS. Notably, FCCN and NOFLITE underperform in certain ACIC 2016 and synthetic settings, likely due to limited training data or suboptimal hyperparameter tuning (Figure 2). Across all experiments, the CCT- and CMC meta-learners consistently strike a balance between calibration and precision, achieving robust CRPS improvements over state-of-the-art alternatives.
@@ A.1 Conformal prediction
[…]
This procedure, termed full or transductive CP, is computationally intensive, as it requires refitting of the underlying model for each $y$ in the target domain. Therefore, a more applicable variant of full CP, called inductive or split CP (ICP) [Papadopoulos et al., 2002], offers computational efficiency and enables the integration of CP with ML algorithms, such as neural networks and tree-based models. In ICP, the training sequence $Z_{1:n}=\{(X_{1},Y_{1}),...,(X_{n},Y_{n})\}$ is portioned into a proper training sequence $\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ and a calibration sequence $\{(X_{m+1},Y_{m+1}),...,(X_{n},Y_{n})\}$. We first train a regression model $\hat{y}(x)$ using the proper training sequence. Then, based on nonconformity score $s$, we generate nonconformity scores $S_{i}$ for data points $(X_{i},Y_{i})$ where $m<i\leq n$. The nonconformity score function $s$ leverages the pre-trained regression model, e.g., the absolute error s=|yy^|s=|y-$\hat${y}|. These nonconformity scores together with test object $X_{n+1}$ and significance level $\alpha$ defines: